Suppose the correlation between height and weight for adults is +0.40. What proportion (or percent) of the variability in weight that is related to height? (Hint-coefficient of determination). (1 pt) please show all work and steps

a. 40%

b. 16%

c. 0.40%

d. 0.16%
Answer #1

We know that the coefficient of determination R2 is given by:

R^2=r^2=(0.40)^2=0.16

So, 16% of the variability in weight that is related to height.

Hence option B.